# Oganesson versus Uranium Hydrogen-like Ions from the Viewpoint of Old Quantum Mechanics

### v3

@@ Abstract.
We compare, within the framework of the old Bohr-Sommerfeld atomic model, Uranium (Z=92)(Z=92)\/ versus hypothetical Oganesson (Z=118)(Z=118)\/ relativistic hydrogen-like ions and beyond. Existence of a self-intersecting orbit in the super strong static Coulomb field is demonstrated with the aid of Mathematica computer algebra system. A possibility of a similar ‘Oganesson-type’ effect in a strong gravitational field is also mentioned. Appearance of a multi self-intersecting trajectory is demonstrated in a hypothetical Unbibium, an element with Z = 122.
@@
Arnold Sommerfeld [30, p. 258]
1. Introduction: Fine Structure Formula and Orbits
@@
As stated in original works [5, 18, 30], classical relativistic Kepler orbits have the form of conic sections as in the nonrelativistic case, but with a new angular variable $\phi=\omega\theta.$ Thus, for elliptical orbits (bound states), the motion from one perihelion ($\phi=0$) to the next ($\phi=2\pi$) requires $\theta=2\pi/\omega\,$, with a per-revolution shift of $\Delta\theta=2\pi/\omega-2\pi.$ [^1] We follow the terminology of classical work [30] where sufficiently small $\Delta\theta$ are discussed.
@@
where $a_{0}=\hbar^{2}/(me^{2})$ is the familiar Bohr radius and $\alpha=e^{2}/(\hbar c)$ is the fine structure formula. These formulas generalize the circular orbits. [^2] Classical solutions of the relativistic Kepler problem are also discussed in [11, pp. 481–482] and [18, pp. 100–102].
2. Some Trajectories of Finite Motion: Uranium vs Oganesson and Beyong
We believe that utilizing the Bohr–Sommerfeld model to contrast $U^{91+}$, $Og^{117+}$, and even hypothetical $Ubb^{121+}$ hydrogen-like ions gives a fresh pedagogical and historical perspective – it links early quantum theory [29, 30] to modern superheavy transuranium relativistic systems [22] in the quasi-classical approximation (using the Dirac equation as a next step [3]).
@@ 2.1. Uranium
In the Bohr-Sommerfeld atomic model under consideration, even for the strong electric field of the Uranium nucleus, a single electron stays on classical relativistic elliptical orbits with advancing perihelion, in a form of open “rosettes” [5, 18, 30] (see Figure 1 with the standard orbit ‘winding number’ one).
Figure 1. Kepler’s Elliptical Counterclockwise Motion in Uranium Hydrogen-like Ion $U^{91+}\/.$
In the above animation of the relativistic Kepler motion of an electron in hydrogen-like Uranium ion, $U^{91+}$, when $Z=92$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.741\ 135$, $\epsilon=0.904\ 882$, and a=0.035 3163a=0.035\ 3163, in Bohr’s atomic units. The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.003\ 3592\ 09$ and $r_{\text{max}}=a(1+\epsilon)=0.067\ 2734\ 72$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, are not shown for simplicity. Counterclockwise rotation of the ellipse over one period is given by $\Delta\theta=2\pi\left(1/\omega-1\right)=$ 2.194 61.2.194\ 61.
@@ 2.2. Oganesson
Our Mathematica animations show that in a super strong electric field of the Oganesson nucleus, the classical relativistic finite Kepler orbits of a single electron may change its topology (see Figure 2 with the orbit ‘winding number’ two [^3] The so-called “chemical orbit” or “double necklace” in terminology of the American cosmologist John Archibald Wheeler [6, 24, 32]. ).
Figure 2. Kepler’s ‘Elliptical’ Clockwise Motion with Self-intersection in Hypothetical Oganesson Hydrogen-like Ion $Og^{117+}\/.$
@@
Figure 3. Kepler’s ‘Elliptical’ Counterclockwise Motion with Four Self-intersections in Hypothetical Unbibium Hydrogen-like Ion $Ubb^{121+}\/.$
@@ 2.3. Unbibium
As another example, let us discuss a hypothetical element with $Z=122$, which has a temporary name, Unbibium (Ubb), one of a yet-to-be-discovered superheavy element at the end of periodic table.
In the above animation of the relativistic Kepler motion of an electron in a hypothetical hydrogen-like Unbibium ion, $Ubb^{121+}$, when $Z=122$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.455\ 419$, $\epsilon=0.949\ 782$, and a=0.020 3534a=0.020\ 3534, in Bohr’s atomic units.
The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.001\ 02211$ and $r_{\text{max}}=a(1+\epsilon)=0.039\ 6847$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, see (see Figure 3, are not shown for simplicity). Rotation of the ellipse over one period is given by $\Delta\theta=2\pi(1/\omega-1)=7.513\ 3.\/$ The ‘winding’ number is three.
@@
Orbit parameters under consideration are compared in the following table for the reader’s convenience:
Parameters: $U^{91+}$ Ion $Og^{117+}$ Ion $Ubb^{121+}$ Ion
$\omega$ 0.741 1350.741\ 135 0.508 4570.508\ 457 0.455 4190.455\ 419
$\epsilon$ 0.904 8820.904\ 882 0.941 4790.941\ 479 0.949 7820.949\ 782
$a/a_{0}$ 0.035 31630.035\ 3163 0.022 20410.022\ 2041 0.020 35340.020\ 3534
$r_{\text{min}}$ 0.003 3592 090.003\ 3592\ 09 0.001 29940.001\ 2994 0.001 022110.001\ 02211
$r_{\text{max}}$ 0.067 2734 720.067\ 2734\ 72 0.043 10870.043\ 1087 0.039 68470.039\ 6847
$\Delta\theta$ 2.194 612.194\ 61 6.074 186.074\ 18 7.513 37.513\ 3
‘Winding’ number 11 22 33
@@ 3. Conclusion
We have demonstrated that in a super strong static electric field, similar to the hypothetical Oganesson hydrogen-like ion, there is a possibility of self-intersecting trajectory, which is different from the standard classical orbits. A possible rotation of an incoming electron around the Oganesson nucleus during the scattering should also be investigated [29]. It would be interesting to analyse a possibility of a similar ‘Oganesson-type’ effect in the theory of gravitation [18] and to study non-standard consequences in quantum electrodynamics in the strong, U, versus super strong, Og, Coulomb fields [1, 9, 10, 12, 13, 14, 19, 22, 23, 27, 28]. For example, in the creation of an electron-positron pair, the electron may be captured by the nucleus directly from the internal orbit. Further nuclear reaction with one of Oganesson’s proton may result in neutron and neutrino thus contributing into the instability of this super heavy element. In a hypothetical Unbibium, element with Z=122,Z=122\/, there is a possibility of a multiple self-intersecting trajectory, which should even more contribute into an orbit electron capture and the corresponding Ubb transformation.
Acknowledgments We are grateful to Dr. Ruben Abagyan, Dr. Sergey I. Kryuchkov, Dr. Eugene Stepanov, and Dr. Alexei Zhedanov for their insightful comments and help.

### v4

# Oganesson versus Uranium Hydrogen-like Ions and Beyond

@@ Abstract.
We compare, within the framework of the old Bohr-Sommerfeld atomic model, Uranium (Z=92)(Z=92)\/ versus hypothetical Oganesson (Z=118)(Z=118)\/ relativistic hydrogen-like ions and beyond. Existence of self-intersecting orbits in the super strong static Coulomb field of superheavy transuranium elements, including the Oganesson, is demonstrated with the aid of Mathematica computer algebra system. A possibility of a similar ‘Oganesson-type’ effect in a strong gravitational field is also mentioned. Appearance of a multi self-intersecting trajectory is demonstrated in a hypothetical Unbibium, an element with Z = 122 and beyond with $Z\leq 135$. A topological classification of the strength for Coulomb fields in the superheavy transuranium elements may be given in terms of the ‘winding’ numbers.
@@
Arnold Sommerfeld [31, p. 258]
1. Transuranics, Fine Structure Formula, and Orbits
Transuranic elements are chemical elements with atomic numbers greater than $92$ (the atomic number of uranium). All of them are unstable and radioactive, and most must be synthesized in a laboratory [17]. The high positive charge of the nucleus creates an extremely strong electric field that affects the behavior of all orbiting electrons. For the innermost electrons, the electrostatic attraction is so strong that their velocities approach the speed of light.
The Bohr-Sommerfeld model was an early attempt to explain atomic structure by incorporating elliptical orbits and special relativity into the Bohr model [3, 16, 18]. However, its simplified, semiclassical approach has fundamental limitations that prevent it from accurately describing transuranic elements. Nonetheless, we would like to explore this “missing opportunity” for highly charged superheavy one-electron systems.
@@ 1. Transuranics, Fine Structure Formula, and Orbits
As stated in original works [5, 19, 31], classical relativistic Kepler orbits have the form of conic sections as in the nonrelativistic case, but with a new angular variable $\phi=\omega\theta.$ Thus, for elliptical orbits (bound states), the motion from one perihelion ($\phi=0$) to the next ($\phi=2\pi$) requires $\theta=2\pi/\omega\,$, with a per-revolution shift of $\Delta\theta=2\pi/\omega-2\pi.$ [^1] In the terminology of classical work [31] where only sufficiently small $\Delta\theta$ are discussed (Figure 1). Throughout this article we shall keep the original name, ‘elliptical’ curve, although for certain $\omega$’s a different topological structure may occur such as orbits with different ‘winding’ numbers and with multiple points of self-intersections (see Figures 2 to 7 below).
@@
where $a_{0}=\hbar^{2}/(me^{2})$ is the familiar Bohr radius and $\alpha=e^{2}/(\hbar c)$ is the fine structure formula. These formulas generalize the circular orbits. [^2] Classical solutions of the relativistic Kepler problem are also discussed in [12, pp. 481–482] and [19, pp. 100–102].
2. Some Trajectories of Finite Motion: Uranium vs Oganesson and Beyond
We believe that utilizing the Bohr–Sommerfeld model to contrast $U^{91+}$, $Og^{117+}$, hypothetical $Ubb^{121+}$ hydrogen-like ions, and beyond gives a fresh pedagogical and historical perspective – it links early quantum theory [30, 31] to modern superheavy transuranium relativistic systems [21] in the quasi-classical approximation (using the Dirac equation as a next step [3]).
@@ 2.1. Uranium
In the Bohr-Sommerfeld atomic model under consideration, even for the strong electric field of the Uranium nucleus, a single electron stays on classical relativistic elliptical orbits with advancing perihelion, in a form of open “rosettes” [5, 19, 31] (see Figure 1 with the standard orbit ‘winding’ number one).
Figure 1. Kepler’s Classical Elliptical Curve Counterclockwise Motion in Relativistic Uranium Hydrogen-like Ion $U^{91+}\/.$ The ‘winding’ number is one.
In the above animation of the relativistic Kepler motion of an electron in hydrogen-like Uranium ion, $U^{91+}$, when $Z=92$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.741\ 135$, $\epsilon=0.904\ 882$, and a=0.035 3163a=0.035\ 3163, in Bohr’s atomic units. The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.003\ 3592\ 09$ and $r_{\text{max}}=a(1+\epsilon)=0.067\ 2734\ 72$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit [31], are not shown for simplicity, Figure 1. Counterclockwise rotation of the ellipse over one period is given by $\Delta\theta=2\pi\left(1/\omega-1\right)=$ 2.194 61.2.194\ 61.
Figure 2. Kepler’s ‘Elliptical’ Curve Clockwise Motion with Self-intersection in Hypothetical Copernicium Hydrogen-like Ion $Cn^{111+}\/.$ The ‘winding’ number is two.
2.2. Copernicium
Copernicium (Cn) is a superheavy, synthetic, and radioactive chemical element with atomic number $Z=112$, discovered in 1996 at the Gesellschaft für Schwerionenforschung (GSI) in Darmstadt, Germany. It was officially named after the astronomer Nicolaus Copernicus in 2010 [17].
Our Mathematica animations show that in a super strong electric field of the Copernicium nucleus, the classical relativistic finite Kepler orbits of a single electron may change its topology, see Figure 2 with the orbit ‘winding’ number two [^3] The so-called “chemical orbit” or “double necklace” in terminology of the American cosmologist John Archibald Wheeler [6, 25, 34]. .
In the above animation of the relativistic Kepler motion of an electron in hydrogen-like Copernicium ion, $Cn^{111+}$, when $Z=112$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.576\ 207$, $\epsilon=0.930\ 786$, and a=0.024 9872a=0.024\ 9872, in Bohr’s atomic units. The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.001\ 7294\ 7$ and $r_{\text{max}}=a(1+\epsilon)=0.048\ 2449$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit are not shown for simplicity. Total rotation of the self-intersecting ‘ellipse’ over one period is given by $\Delta\theta=2\pi\left(1/\omega-1\right)=4.621\ 2.$
[…]
@@ 2.3. Oganesson
Our Mathematica animations show that in a super strong electric field of the Oganesson nucleus, the classical relativistic finite Kepler orbits of a single electron may change its topology, see Figure 3 with the orbit ‘winding’ number two
Figure 3. Kepler’s ‘Elliptical’ Curve Clockwise Motion with Self-intersection in Hypothetical Oganesson Hydrogen-like Ion $Og^{117+}\/.$ The ‘winding’ number is two.
@@
Figure 4. Kepler’s ‘Elliptical’ Counterclockwise Motion with Four Self-intersections in Hypothetical Unbibium Hydrogen-like Ion $Ubb^{121+}\/.$ The ‘winding’ number is three.
@@ 2.4. Unbibium
Element $122$, is a hypothetical chemical element with the placeholder symbol $Ubb$. In the above animation of the relativistic Kepler motion of an electron in a hydrogen-like Unbibium ion, $Ubb^{121+}$, the atomic number $Z=122$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.455\ 419$, $\epsilon=0.949\ 782$, and a=0.020 3534a=0.020\ 3534, in Bohr’s atomic units.
The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.001\ 02211$ and $r_{\text{max}}=a(1+\epsilon)=0.039\ 6847$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, see Figure 4, are not shown for simplicity. Rotation of the ellipse over one period is given by $\Delta\theta=2\pi(1/\omega-1)=7.513\ 3.\/$ The ‘winding’ number is three.
@@
Figure 5. Kepler’s ‘Elliptical’ Curve Counterclockwise Motion with Four Self-intersections in Hypothetical Unbiennium Hydrogen-like Ion $Ube^{128+}\/.$ The ‘winding’ number is four.
Figure 6. Kepler’s ‘Elliptical’ Curve Counterclockwise Motion with Multiple Self-intersections in Hypothetical Untribium Hydrogen-like Ion $Utb^{131+}\/.$ The ‘winding’ number is five.
Figure 7. Kepler’s ‘Elliptical’ Curve Counterclockwise Motion with Multiple Self-intersections in Hypothetical Unbiennium Hydrogen-like Ion $Utp^{134+}\/.$ The ‘winding’ number is six.
2.5. Unbiennium
In the above animation of the relativistic Kepler motion of an electron in a hypothetical hydrogen-like Unbiennium ion, $Ube^{128+}$, when $Z=129$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.337\ 408$, $\epsilon=0.967\ 653$, and a=0.016 9559a=0.016\ 9559, in Bohr’s atomic units.
Once again, the perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.000\ 5484\ 74$ and $r_{\text{max}}=a(1+\epsilon)=0.033\ 3634$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, are not shown for simplicity (Figure 4). Rotation of the ellipse over one period is given by $\Delta\theta=2\pi(1/\omega-1)=12.33\ 87.\/$ The ‘winding’ number is four.
2.6. Untribium
In the above animation of the relativistic Kepler motion of an electron in a hypothetical hydrogen-like Unbiennium ion, $Utb^{131+}$, when $Z=132$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.268\ 605$, $\epsilon=0.977\ 328$, and a=0.015 3084a=0.015\ 3084, in Bohr’s atomic units.
The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.000\ 3470\ 77$ and $r_{\text{max}}=a(1+\epsilon)=0.030\ 2698$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, see Figure 6, are not shown for simplicity. Rotation of the ellipse over one period is given by $\Delta\theta=2\pi(1/\omega-1)=17.10\ 88.\/$ The ‘winding’ number is five.
2.7. Untripentium
In the above animation of the relativistic Kepler motion of an electron in a hypothetical hydrogen-like Unbiennium ion, $Utp^{134+}$, when $Z=135$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.171\ 738$, $\epsilon=0.989\ 201$, and a=0.013 287a=0.013\ 287, in Bohr’s atomic units.
The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.000\ 14349$ and $r_{\text{max}}=a(1+\epsilon)=0.026\ 4305$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, see Figure 7, are not shown for simplicity. Rotation of the ellipse over one period is given by $\Delta\theta=2\pi(1/\omega-1)=30.30\ 26.\/$ The ‘winding’ number is six.
Orbit parameters under consideration are compared in the following tables for the reader’s convenience:
$Z$ $U^{91+}$ Ion $Cn^{111+}$ Ion $Og^{117+}$ Ion
$\omega$ 0.741 1350.741\ 135 0.576 2070.576\ 207 0.508 4570.508\ 457
$\epsilon$ 0.904 8820.904\ 882 0.930 7860.930\ 786 0.941 4790.941\ 479
$a/a_{0}$ 0.035 31630.035\ 3163 0.024 98720.024\ 9872 0.022 20410.022\ 2041
$r_{\text{min}}$ 0.003 3592 090.003\ 3592\ 09 0.001 729470.001\ 72947 0.001 29940.001\ 2994
$r_{\text{max}}$ 0.067 2734 720.067\ 2734\ 72 0.048 24490.048\ 2449 0.043 10870.043\ 1087
$\Delta\theta$ 2.194 612.194\ 61 4.621 24.621\ 2 6.074 186.074\ 18
[…]
$Z$ $Ubb^{121+}$ Ion $Ube^{128+}$ Ion $Utb^{131+}$ Ion $Utp^{134+}$ Ion
$\omega$ 0.455 4190.455\ 419 0.337 4080.337\ 408 0.268 6050.268\ 605 0.171 7380.171\ 738
$\epsilon$ 0.949 7820.949\ 782 0.967 6530.967\ 653 0.977 3280.977\ 328 0.989 2010.989\ 201
$a/a_{0}$ 0.020 35340.020\ 3534 0.016 95590.016\ 9559 0.015 30840.015\ 3084 0.013 2870.013\ 287
$r_{\text{min}}$ 0.001 022110.001\ 02211 0.000 5484 740.000\ 5484\ 74 0.000 3470 770.000\ 3470\ 77 0.000 143490.000\ 14349
$r_{\text{max}}$ 0.039 68470.039\ 6847 0.033 36340.033\ 3634 0.030 26980.030\ 2698 0.026 43050.026\ 4305
$\Delta\theta$ 7.513 37.513\ 3 12.33 8712.33\ 87 17.10 8817.10\ 88 30.30 2630.30\ 26
[…]
@@ 3. Conclusion
We have demonstrated that in a super strong static electric field, similar to the hypothetical Oganesson hydrogen-like ion, there is a possibility of self-intersecting trajectory, which is different from the standard classical orbits. A possible rotation of an incoming electron around the Oganesson nucleus during the scattering should also be investigated [30]. It would be interesting to analyze a possibility of a similar ‘Oganesson-type’ effect in the theory of gravitation [19] and to study non-standard consequences in quantum electrodynamics in the strong, U, versus super strong, Og and beyond, Coulomb fields [1, 10, 11, 13, 14, 15, 20, 21, 24, 28, 29]. For example, in the creation of an electron-positron pair, the electron may be captured by the nucleus directly from the internal orbit. Further nuclear reaction with one of Oganesson’s proton may result in neutron and neutrino, thus contributing into the instability of this super heavy element. In a hypothetical Unbibium, element with Z=122,Z=122\/, there is a possibility of a multiple self-intersecting trajectory, which should even more contribute into an orbit electron capture and the corresponding Ubb transformation and beyond, when $Z\leq 135\/$.
Our examples may give an idea of a topological classification, – by orbits ‘winding’ numbers –, for the strength of Coulomb fields in the superheavy transuranium elements.
Sommerfeld vs oganesson puzzles: As well-known, the energy levels of a hydrogen-like system in Dirac’s theory match, precisely, the Sommerfeld formula (1.1) – an outcome known as the “Sommerfeld Puzzle” [5], discussed further in [8, pp. 426–429]. This “puzzle” has recently been resolved [33]. But why does the presence of an extra loop in the ‘old’ Bohr-Sommerfeld, or semi-classical, trajectory of the electron in the transuranic heavy ions under consideration coincide with the parameters of artificially created elements on Earth, including Oganesson, which is located at the end of the Mendeleev periodic table?
Acknowledgments We are grateful to Professor Ruben Abagyan, Professor Victor V. Dodonov, Dr. Sergey I. Kryuchkov, Dr. Eugene Stepanov, and Professor Alexei Zhedanov for their insightful comments and help.